# WHAM: Reconstructing World-grounded Humans with Accurate 3D Motion

### v1

@@ Abstract
The estimation of 3D human motion from video has progressed rapidly but current methods still have several key limitations. First, most methods estimate the human in camera coordinates. Second, prior work on estimating humans in global coordinates often assumes a flat ground plane and produces foot sliding. Third, the most accurate methods rely on computationally expensive optimization pipelines, limiting their use to offline applications. Finally, existing video-based methods are surprisingly less accurate than single-frame methods. We address these limitations with WHAM (World-grounded Humans with Accurate Motion), which accurately and efficiently reconstructs 3D human motion in a global coordinate system from video. WHAM learns to lift 2D keypoint sequences to 3D using motion capture data and fuses this with video features, integrating motion context and visual information. WHAM exploits camera angular velocity estimated from a SLAM method together with human motion to estimate the body’s global trajectory. We combine this with a contact-aware trajectory refinement method that lets WHAM capture human motion in diverse conditions, such as climbing stairs. WHAM outperforms all existing 3D human motion recovery methods across multiple in-the-wild benchmarks. Code will be available for research purposes at http://wham.is.tue.mpg.de/.
@@ 1 Introduction
Our goal is to accurately estimate the 3D pose and shape of a person from monocular video. This is a longstanding problem and, while the field has made rapid progress, several key challenges remain. First, human motion should be computed in a consistent global coordinate system. Second, the method should be computationally efficient, supporting real-time processing. Third, the results should be accurate, temporally smooth, detailed, natural looking, and have realistic foot-ground contact. Fourth, the capture should work with an arbitrary moving camera. These constraints need to be satisfied to make markerless human motion capture widely available for applications in gaming, AR/VR, autonomous driving, sports analysis, and human-robot interaction. We address these challenges with WHAM (World-grounded Humans with Accurate Motion), which enables fast and accurate recovery of 3D human motion from a moving camera; see Fig. WHAM: Reconstructing World-grounded Humans with Accurate 3D Motion.
@@
It seems natural that, in estimating 3D humans from video, we should be able to exploit the temporal nature of video. Counter-intuitively, existing video-based methods for 3D human pose and shape (HPS) estimation [15, 17, 30, 6, 51, 43] are less accurate than the best single-frame methods [14, 21, 18, 59, 23, 59, 25, 7]. This may be an issue of training data. There are large datasets of single images with ground-truth 3D human poses containing a diversity of body shapes, poses, backgrounds, lighting, etc. In contrast, video datasets with ground truth are much more limited.
@@
To address this, WHAM leverages both the large-scale AMASS motion capture (mocap) dataset [32] and video datasets. Our key idea is to learn about 3D human motion from AMASS and then learn to fuse this information with temporal image cues from video, getting the best of both. Similar to previous work [61, 60], we use AMASS to generate synthetic 2D keypoints and ground-truth motion sequence pairs, from which we pretrain a motion encoder and decoder to lift sequences of 2D keypoints to sequences of 3D poses. This motion encoder captures the essential motion context recursively over time, while the decoder translates this context into realistic 3D motion. Given the robustness of recent 2D keypoint detection models [53, 58], our pre-trained model does a good job of predicting human pose from video.
@@
Keypoints alone, however, are too sparse for accurate, unambiguous, 3D mesh estimation. To improve accuracy, we jointly train a new feature integrator network that merges information from video sequences and 2D-keypoint sequences. Like the motion context, the image features are integrated over time and the image encoder and feature integrator are trained using video datasets with known 3D pose and shape [49, 11, 33, 15]. This integration process supplements the motion context extracted from the sparse input signal (i.e., 2D keypoints) with dense visual context, significantly improving the recovered pose and shape accuracy.
@@
While the above approach produces accurate motion, we want this motion in a global coordinate. Most previous methods, compute the body in camera coordinates. Estimating the global human trajectory is particularly challenging when the camera is moving because the motions of the body and the camera are entangled. Recent work addresses this challenge by exploiting a learned prior distribution over human motions, together with camera information from SLAM methods [54, 41, 20] or dense 3D scene information from COLMAP [28], to solve for global human motion using optimization. However, these optimization-based methods are computationally expensive and far from real time. Recent regression-based methods are faster but either constrain the problem with static or known camera conditions [60, 42] or have temporal jitter and limited accuracy [45]. We tackle this challenge with two additional modules. First, we predict the global orientation and egocentric velocity of the human from the motion context by training a global trajectory decoder. Specifically, we concatenate the camera’s angular velocity to the context and train the global trajectory decoder to recursively predict the current orientation and root velocity, effectively factoring camera motion from human motion. Global translation is computed by roll-out. WHAM takes the camera’s angular velocity either from the output of a SLAM method or from a camera’s gyroscope when available. We demonstrate the use of both.
@@
The above solution relies on knowledge of human motion learned from AMASS. While this works well when people are moving on a flat ground plane, it can fail to capture elevation changes when the surface is not flat, e.g. when ascending the stairs. Even though we have no explicit ground-plane assumption (unlike [54]), this occurs because AMASS has a limited amount of such data. To address this, we introduce foot contact as an additional explicit form of motion context. We train WHAM to predict the likelihood of foot-ground contact using estimated contact labels from both the AMASS and 3D video datasets. We then train a trajectory refinement network that outputs an update to the root orientation and velocity based on the information about the foot contact/velocity. This refinement based on foot contact enables WHAM to accurately estimate human motion in a global coordinate system even when the terrain is not flat.
@@
WHAM has very low computational overhead because it is an on-line algorithm that recursively predicts the pose, shape, and global motion parameters. The network, excluding preprocessing (bounding box detection, keypoint detection, and person identification), runs at 200 fps, significantly faster than prior methods. Also, despite not using global optimization like [54], we obtain accurate 3D camera trajectories and global body motions with minimal drift. Through extensive comparisons on multiple in-the-wild datasets as well as detailed ablation studies, we find that WHAM achieves state-of-the-art (SOTA) accuracy on 3D human pose estimation as well as global trajectory estimation (see Fig. WHAM: Reconstructing World-grounded Humans with Accurate 3D Motion).
@@
In summary, in this paper we: (1) introduce the first approach to effectively fuse 3D human motion context and video context for 3D HPS regression; (2) propose a novel global trajectory estimation framework that leverages motion context and foot contact to effectively address foot sliding and enable the 3D tracking of people on non-planar surfaces; (3) perform HPS in global coordinates efficiently; (4) achieve state-of-the-art (SOTA) performance on multiple in-the-wild benchmark datasets (3DPW [49], RICH [10], EMDB [16]). WHAM is the first video-based method to outperform all image-based and video-based methods on per-frame accuracy (MPJPE, PA-MPJPE, and PVE) while maintaining temporal smoothness (Accel). Pretrained models and training code will be available for research purposes.
@@ 2 Related Work
Many existing methods follow the architecture of HMR [14], which uses a pre-trained backbone to predict image features followed by a multilayer perceptron (MLP) that regresses SMPL [29] pose parameters from image features. To train such networks it leverages paired images with SMPL parameters, which are often pseudo-groundtruth SMPL parameters estimated from 2D keypoints [21, 12, 35, 36, 25]. Other architectures for HPS regression have also been proposed [18, 23, 59, 24, 7, 19]. None of these methods use video or estimate the body in global coordinates. While quite accurate, when these image-based models are applied independently to frames of a video sequence, the shape and pose can be temporally inconsistent. In contrast, WHAM effectively aggregates temporal information to provide frame-accurate and temporally-coherent 3D HPS estimation.
@@
Given limited video training data with ground truth SMPL poses, several single-frame methods infer a mesh from 2D/3D keypoints [5, 8, 39, 31, 34] and use the keypoints as a proxy for training. Another set of approaches exploits 3D mocap data, which is plentiful [32], to train a network to lift 2D joints to 3D poses, which are used as a proxy for ground truth 3D. MotionBERT [61] synthesizes 2D keypoints through orthographic projection to learn unified motion representation. ProxyCap [60] creates virtual cameras with heuristic pose distribution, on which synthetic 3D keypoints are projected. Despite benefiting from the scale of the mocap dataset, these approaches do not fully utilize the visual information available in the video at test time. In our work, we propose a combined network architecture and training strategy that leverages both proxy representations of human pose (lifting) and visual context extracted from video.
@@
Global 3D Motion Estimation with Dense Sensors. Several methods augment video data with other sensors to estimate 3D HPS in world coordinates. The 3DPW dataset [49] employs pre-calibrated body-worn inertial sensors and a handheld camera to jointly optimize the camera and human motion in challenging environments. Similarly, the EMDB dataset [16] uses electromagnetic sensors with an RGB-D camera, enabling accurate human motion capture in the world. Although body-worn sensors aid reconstruction of global human motion, they are intrusive, require cooperation, and do not help with archival video. BodySLAM++ [9] introduces a rapid optimization method using a visual-inertial sensor, comprising stereo cameras and an IMU. In contrast, we use a standard monocular camera, balancing accessibility and accuracy without the need for specialized equipment. While WHAM can take the camera gyro as input, this is not required.
@@
Monocular Global 3D Human Trajectory Estimation. Estimating the global human trajectory from a monocular dynamic camera is challenging. Previous work relies on learned prior distributions of human motion to separate human motion from camera motion. GLAMR [57] predicts the global trajectory based on a predicted and infilled 3D motion sequence and optimizes it across multiple individuals in the scene. However, since GLAMR does not consider camera motion cues, the output trajectory may be noisy when the camera is rotating. SLAHMR [54] and PACE [20] use off-the-shelf SLAM algorithms [46, 47] and jointly optimize the camera and human motion to minimize the negative log likelihood of a learned motion prior [40]. While they achieve good results, their optimization approach is computationally expensive. TRACE [45] is a pure regression method that utilizes optical flow as a motion cue and estimates multiple people at once, but lacks temporal consistency. GloPro [42] regresses the uncertainty of the global human motion in real time, but requires known camera poses. In contrast, WHAM leverages both explicit and implicit prior knowledge of human motion and efficiently reconstructs accurate and temporally coherent 3D human motion in world coordinates.
@@
Figure 2: An Overview of WHAM. WHAM takes the sequence of 2D keypoints estimated by a pretrained detector and encodes it to the motion feature. WHAM then updates the motion feature using another sequence of image features extracted from the image encoder through the feature integrator. From the updated motion feature, the Local Motion Decoder estimates 3D motion in the camera coordinate system and foot-ground contact probability. The Trajectory Decoder takes the motion feature and camera angular velocity to initially estimate the global root orientation and egocentric velocity, which are then updated through the Trajectory Refiner using the foot-ground contact. The final output of WHAM is pixel-aligned 3D human motion with the 3D trajectory in the global coordinates.
@@ 3.2 Network Architecture
Uni-directional Motion Encoder and Decoder. In contrast to existing methods [6, 30, 51, 43, 61], which use windows with a fixed time duration, we use uni-directional recurrent neural networks (RNN) for the motion encoder and decoder, making WHAM suitable for online inference. The objective of the motion encoder $E_{M}$ is to extract the motion context $\phi_{m}^{(t)}$ from the current and previous sequence of 2D keypoints and the initial hidden state $h_{E}^{(0)}$:
@@
We normalize keypoints to a bounding box around the person and concatenate the box’s center and scale to the keypoints, similar to CLIFF [25]. The role of the motion decoder $D_{M}$ is to recover SMPL parameters $(\theta,\beta)$, weak-perspective camera translation $c$, and foot-ground contact probability $p$, from the motion feature history:
@@
Here, $\hat{\phi}_{m}^{(t)}$ is the motion feature integrated with the image feature $\phi_{i}^{(t)}$ (described below). During pre-training on synthetic data, the image feature is not available and we set $\hat{\phi}_{m}^{(t)}=\phi_{m}^{(t)}$. As the encoder and decoder are tasked with reconstructing a dense 3D representation $\Theta$ from a sparse 2D input signal $x_{2D}$, we design an intermediate task to predict the 3D keypoints $x_{3D}$ and use them as the intermediate motion representation. This cascaded approach guides $\phi_{m}$ to represent the implicit context of motion and the 3D spatial structure of the body. We initialize the hidden states of the motion encoder and decoder, $\big{(}h_{E}^{(0)},h_{D}^{(0)}\big{)}$, following PIP [56]; See Sup. Mat. for details.
@@
This supplements motion features pre-trained on the 2D-to-3D lifting task using AMASS with visual context, resulting in enriched motion features that use image evidence to help disambiguate the task.
Global Trajectory Decoder. We design an additional decoder, $D_{T}$, to predict the rough global root orientation $\Gamma_{0}^{(t)}$ and root velocity $v_{0}^{(t)}$ from the motion feature $\phi_{m}^{(t)}$. Since $\phi_{m}$ is derived from the input signals in the camera coordinates, it is highly challenging to decouple the human and camera motion from it. To address this ambiguity, we append the angular velocity of the camera, $\omega^{(t)}$, to the motion feature, $\phi_{m}^{(t)}$, to create a camera-agnostic motion context. This design choice makes WHAM compatible with both off-the-shelf SLAM algorithms [47, 46] and gyroscope measurements that are widely available from modern digital cameras. We recursively predict global orientation, $\Gamma_{0}^{(t)}$, using the uni-directional RNN.Similar to GLAMR [57], we use the ego-centric root velocity to make the prediction invariant to the global orientation:
@@
Contact Aware Trajectory Refinement. Good 3D motion in world coordinates in most scenarios implies accurate foot-ground contact without sliding. We want WHAM to generalize beyond flat ground planes, which are typically assumed in prior work. Specifically, our new trajectory refiner aims to resolve foot sliding and enables WHAM to generalize well to diverse motions, including climbing stairs. The refinement involves two stages. First, we adjust the ego-centric root velocity to $\tilde{v}^{(t)}$ to minimize foot sliding, based on the foot-ground contact probability $p^{(t)}$:
@@
where $\bar{v}_{f}^{(t)}$ is the averaged velocity of the toes and heels in the world coordinate when their contact probability, $p^{(t)}$, is higher than a threshold. However, this velocity adjustment often introduces noisy translation when the contact and pose estimation is inaccurate. Therefore, we propose a simple learning mechanism in which a trajectory refining network, $R_{T}$, updates the root orientation and velocity to address this issue. Finally, the global translation is computed through a roll-out operation:

### v2

@@ Abstract
The estimation of 3D human motion from video has progressed rapidly but current methods still have several key limitations. First, most methods estimate the human in camera coordinates. Second, prior work on estimating humans in global coordinates often assumes a flat ground plane and produces foot sliding. Third, the most accurate methods rely on computationally expensive optimization pipelines, limiting their use to offline applications. Finally, existing video-based methods are surprisingly less accurate than single-frame methods. We address these limitations with WHAM (World-grounded Humans with Accurate Motion), which accurately and efficiently reconstructs 3D human motion in a global coordinate system from video. WHAM learns to lift 2D keypoint sequences to 3D using motion capture data and fuses this with video features, integrating motion context and visual information. WHAM exploits camera angular velocity estimated from a SLAM method together with human motion to estimate the body’s global trajectory. We combine this with a contact-aware trajectory refinement method that lets WHAM capture human motion in diverse conditions, such as climbing stairs. WHAM outperforms all existing 3D human motion recovery methods across multiple in-the-wild benchmarks. Code is available for research purposes at http://wham.is.tue.mpg.de/.
@@ 1 Introduction
Our goal is to accurately estimate the 3D pose and shape of a person from monocular video. This is a longstanding problem and, while the field has made rapid progress, several key challenges remain. First, human motion should be computed in a consistent global coordinate system. Second, the method should be computationally efficient. Third, the results should be accurate, temporally smooth, detailed, natural looking, and have realistic foot-ground contact. Fourth, the capture should work with an arbitrary moving camera. These constraints need to be satisfied to make markerless human motion capture widely available for applications in gaming, AR/VR, autonomous driving, sports analysis, and human-robot interaction. We address these challenges with WHAM (World-grounded Humans with Accurate Motion), which enables fast and accurate recovery of 3D human motion from a moving camera; see Fig. 1.
@@
It seems natural that, in estimating 3D humans from video, we should be able to exploit the temporal nature of video. Counter-intuitively, existing video-based methods for 3D human pose and shape (HPS) estimation [15, 17, 30, 6, 52, 43] are less accurate than the best single-frame methods [14, 21, 18, 60, 23, 60, 25, 26, 7]. This may be an issue of training data. There are large datasets of single images with ground-truth 3D human poses containing a diversity of body shapes, poses, backgrounds, lighting, etc. In contrast, video datasets with ground truth are much more limited.
@@
To address this, WHAM leverages both the large-scale AMASS motion capture (mocap) dataset [32] and video datasets. Our key idea is to learn about 3D human motion from AMASS and then learn to fuse this information with temporal image cues from video, getting the best of both. Similar to previous work [62, 61], we use AMASS to generate synthetic 2D keypoints and ground-truth motion sequence pairs, from which we pretrain a motion encoder, which captures the motion context, and decoder that lifts sequences of 2D keypoints to sequences of 3D poses. Given the robustness of recent 2D keypoint detection models [54, 59], our pretrained model does a good job of predicting human pose from video.
@@
Keypoints alone, however, are too sparse for accurate 3D mesh estimation. To improve accuracy, we jointly train a feature integrator network that merges information from video and 2D-keypoint sequences. We use a pretrained image encoder from previous work [21, 7, 25, 1] and train the feature integrator using video datasets [49, 11, 15, 33]. This integration process supplements the motion context extracted from the sparse 2D keypoints with dense visual context, significantly improving the recovered pose and shape accuracy.
@@
While the above approach produces accurate motion, we want this motion in global coordinates, unlike most previous methods that compute the body in camera coordinates. Estimating the global human trajectory is challenging when the camera is moving because the motions of the body and the camera are entangled. Recent work addresses this with optimization fitting based on a learned human motion prior and camera information from SLAM methods [55, 41, 20] or dense 3D scene information from COLMAP [28]. However, these methods are computationally expensive and far from real time. Recent regression-based methods are faster but either constrain the problem with static or known camera conditions [61, 42] or have temporal jitter and limited accuracy [45]. We tackle this challenge with two additional modules. First, we predict the global orientation and root velocity of the human from the sequence of 2D keypoints by training a global trajectory decoder. Specifically, we concatenate the camera’s angular velocity to the context and train the global trajectory decoder to recursively predict the current orientation and root velocity, effectively factoring camera motion from human motion. WHAM takes the camera’s angular velocity either from the output of a SLAM method or from a camera’s gyroscope when available.
@@
The above solution relies on knowledge of human motion learned from AMASS. Therefore, it can fail to capture elevation changes when the surface is not flat, e.g. when ascending the stairs because AMASS has a limited amount of such data. To address this, we introduce foot contact as an additional explicit form of motion context. We train WHAM to predict the likelihood of foot-ground contact using estimated contact labels from both AMASS and 3D video datasets. We then train a trajectory refinement network that outputs an update to the root orientation and velocity based on the information about the foot contact/velocity. This refinement enables WHAM to accurately estimate human motion in a global coordinate system even when the terrain is not flat.
@@
WHAM has very low computational overhead because it is an on-line algorithm that recursively predicts the pose, shape, and global motion parameters. The network, excluding preprocessing (bounding box detection, keypoint detection, and person identification), runs at 200 fps, significantly faster than prior methods. Also, despite not using global optimization like [55], we obtain accurate 3D camera trajectories and global body motions with minimal drift. Through extensive comparisons on multiple in-the-wild datasets as well as detailed ablation studies, we find that WHAM achieves state-of-the-art (SOTA) accuracy on 3D human pose estimation as well as global trajectory estimation (see Fig. 1).
@@
In summary, in this paper we: (1) introduce the first approach to effectively fuse 3D human motion context and video context for 3D HPS regression; (2) propose a novel global trajectory estimation framework that leverages motion context and foot contact to effectively address foot sliding and enable the 3D tracking of people on non-planar surfaces; (3) efficiently perform HPS regression in global coordinates; (4) achieve state-of-the-art (SOTA) performance on multiple in-the-wild benchmark datasets (3DPW [49], RICH [10], EMDB [16]). WHAM is the first video-based method to outperform all image-based and video-based methods on per-frame accuracy while maintaining temporal smoothness.
@@ 2 Related Work
Many existing methods follow the architecture of HMR [14], which uses a pretrained backbone to predict image features followed by a multilayer perceptron (MLP) that regresses SMPL [29] pose parameters from image features. Training such networks typically leverages paired images with SMPL parameters; these parameters are often pseudo-groundtruth (p-GT), estimated from 2D keypoints [21, 12, 35, 36, 25]. Other architectures for HPS regression have also been proposed [18, 23, 60, 24, 7, 19]. None of these methods use video or estimate the body in global coordinates. While quite accurate, when these image-based models are applied independently to frames of a video sequence, the shape and pose can be temporally inconsistent. In contrast, WHAM effectively aggregates temporal information to provide frame-accurate and temporally-coherent 3D HPS estimation.
@@
Given limited video training data with ground truth SMPL poses, several single-frame methods infer a mesh from 2D/3D keypoints [5, 8, 39, 31, 34] and use the keypoints as a proxy for training. Another set of approaches exploits 3D mocap data, which is plentiful [32], to train a network to lift 2D joints to 3D poses, which are used as a proxy for ground truth 3D. MotionBERT [62] synthesizes 2D keypoints through orthographic projection to learn a unified motion representation. ProxyCap [61] projects synthetic 3D keypoints into virtual cameras using a heuristic camera pose distribution. Despite benefiting from the scale of mocap datasets, these approaches do not fully utilize the visual information available in the video at test time. Here, we propose a combined network architecture and training strategy that leverages both proxy representations of human pose (lifting) and visual context extracted from video.
@@
Global 3D Motion Estimation with Dense Sensors. Several methods augment video data with other sensors to estimate 3D HPS in world coordinates. The 3DPW dataset [49] employs pre-calibrated body-worn inertial sensors and a handheld camera to jointly optimize the camera and human motion in challenging environments. Similarly, the EMDB dataset [16] uses electromagnetic sensors with an RGB-D camera, enabling accurate human motion capture in the world. While body-worn sensors aid global human motion reconstruction, they are intrusive, require cooperation, and do not help with archival video. BodySLAM++ [9] uses an optimization method with a visual-inertial sensor, comprising stereo cameras and an IMU. In contrast, we use a standard monocular camera, balancing accessibility and accuracy without using specialized equipment. While WHAM can take the camera gyro as input, this is not required.
@@
Monocular Global 3D Human Trajectory Estimation. Estimating the global human trajectory from a monocular dynamic camera is challenging. Previous work relies on learned prior distributions of human motion to separate human motion from camera motion. GLAMR [58] computes the global trajectory based on a predicted and infilled 3D motion sequence and optimizes it across multiple individuals in the scene. However, since GLAMR does not consider camera motion cues, the output trajectory may be noisy when the camera is rotating. SLAHMR [55] and PACE [20] use off-the-shelf SLAM algorithms [46, 47] and jointly optimize the camera and human motion to minimize the negative log-likelihood of a learned motion prior [40]. While they achieve good results, their optimization approach is computationally expensive. TRACE [45] is a pure regression method that utilizes optical flow as a motion cue and estimates multiple people at once, but lacks temporal consistency. GloPro [42] regresses the uncertainty of the global human motion in real-time, but requires known camera poses. In contrast, WHAM leverages both explicit and implicit prior knowledge of human motion and efficiently reconstructs accurate and temporally coherent 3D human motion in world coordinates.
@@
Figure 2: An Overview of WHAM. WHAM takes the sequence of 2D keypoints estimated by a pretrained detector and encodes it into a motion feature. WHAM then updates the motion feature using another sequence of image features extracted from the image encoder through the feature integrator. From the updated motion feature, the Local Motion Decoder estimates 3D motion in the camera coordinate system and foot-ground contact probability. The Trajectory Decoder takes the motion feature and camera angular velocity to initially estimate the global root orientation and egocentric velocity, which are then updated through the Trajectory Refiner using the foot-ground contact. The final output of WHAM is pixel-aligned 3D human motion with the 3D trajectory in the global coordinates.
@@ 3.2 Network Architecture
Uni-directional Motion Encoder and Decoder. In contrast to existing methods [6, 30, 52, 43, 62], which use windows with a fixed time duration, we use uni-directional recurrent neural networks (RNN) for the motion encoder and decoder, making WHAM suitable for online inference. The objective of the motion encoder, $E_{M}$, is to extract the motion context, $\phi_{m}^{(t)}$, from the current and previous sequence of 2D keypoints and the initial hidden state, $h_{E}^{(0)}$:
@@
We normalize keypoints to a bounding box around the person and concatenate the box’s center and scale to the keypoints, similar to CLIFF [25]. The role of the motion decoder, $D_{M}$, is to recover SMPL parameters, $(\theta,\beta)$, weak-perspective camera translation, $c$, and foot-ground contact probability, $p$, from the motion feature history:
@@
Here, $\hat{\phi}_{m}^{(t)}$ is the motion feature integrated with the image feature $\phi_{i}^{(t)}$ (described below). During pretraining on synthetic data, the image feature is not available and we set $\hat{\phi}_{m}^{(t)}=\phi_{m}^{(t)}$. As the encoder and decoder are tasked with reconstructing a dense 3D representation $\Theta$ from a sparse 2D input signal $x_{2D}$, we design an intermediate task to predict the 3D keypoints $x_{3D}$ and use them as the intermediate motion representation. This cascaded approach guides $\phi_{m}$ to represent the implicit context of motion and the 3D spatial structure of the body. Similar to PIP [57], we use Neural Initialization that uses MLP to initialize the hidden state of the motion encoder and decoder, $\big{(}h_{E}^{(0)},h_{D}^{(0)}\big{)}$; see section 4.3 and Sup. Mat. for details.
@@
This supplements motion features pretrained on the 2D-to-3D lifting task using AMASS with visual context, resulting in enriched motion features that use image evidence to help disambiguate the task.
Global Trajectory Decoder. We design an additional decoder, $D_{T}$, to predict the rough global root orientation $\Gamma_{0}^{(t)}$ and root velocity $v_{0}^{(t)}$ from the motion feature $\phi_{m}^{(t)}$. Since $\phi_{m}$ is derived from the input signals in the camera coordinates, it is highly challenging to decouple the human and camera motion from it. To address this ambiguity, we append the angular velocity of the camera, $\omega^{(t)}$, to the motion feature, $\phi_{m}^{(t)}$, to create a camera-agnostic motion context. This design choice makes WHAM compatible with both off-the-shelf SLAM algorithms [47, 46] and gyroscope measurements that are widely available from modern digital cameras. We recursively predict global orientation, $\Gamma_{0}^{(t)}$, using the uni-directional RNN.
@@
Contact Aware Trajectory Refinement. Good 3D motion in world coordinates in most scenarios implies accurate foot-ground contact without sliding. We want WHAM to generalize beyond flat ground planes, which are typically assumed in prior work. Specifically, our new trajectory refiner aims to resolve foot sliding and enables WHAM to generalize well to diverse motions, including climbing stairs. The refinement involves two stages. First, we adjust the ego-centric root velocity to $\tilde{v}^{(t)}$ to minimize foot sliding, based on the foot-ground contact probability $p^{(t)}$, estimated from the motion decoder $D_{M}$:
@@
where $\bar{v}_{f}^{(t)}$ is the averaged velocity of the toes and heels in world coordinates when their contact probability, $p^{(t)}$, is higher than a threshold. However, this velocity adjustment often introduces noisy translation when the contact and pose estimation is inaccurate. Therefore, we propose a simple learning mechanism in which a trajectory refining network, $R_{T}$, updates the root orientation and velocity to address this issue. Finally, the global translation is computed through a roll-out operation: